# Long-Term Multi-Objective Optimization for Integrated Unit Commitment and Investment Planning for District Heating Networks

### v1

@@ Abstract
The need to decarbonize the energy system has intensified the focus on district heating networks in urban and suburban areas. Therefore, exploring transformation pathways with reasonable trade-offs between economic viability and environmental goals became necessary. We introduce a network-flow-based model class integrating unit commitment and long-term investment planning for multi-energy systems. While the integration of unit commitment and investment planning has been applied to multi-energy systems, a formal introduction and suitability for the application of long-term portfolio planning of an energy provider on an urban scale has yet to be met. Based on mixed integer linear programming, the model bridges the gap between overly detailed industrial modeling tools not designed for computational efficiency at scale and rather abstract academic models. The formulation is tested on Berlin’s district heating network. Hence, the challenge lies in a large number of variables and constraints and the coupling of time steps, for example, through investment decisions. A case study explores different solutions on the Pareto front defined by optimal trade-offs between minimizing costs and CO2 emissions through a lexicographic optimization approach. The resulting solution catalog can provide decision-makers valuable insights into feasible transformation pathways, highlighting distinctions between robust and target-dependent investments.
@@ 1 Introduction
While energy providers are urged to decarbonize, current optimization tools for multi-energy systems still fall into two categories: On one hand, detailed industrial modeling tools (e.g,. BoFiT by Volue [^4] https://www.volue.com/products-and-services/volue-bofit-optimisation-chp-sector-coupling-and-industry, ResOpt by Kisters [^5] https://www.kisters.eu/product/forecasting-and-optimisation/) do not allow for an integrated structural optimization for large input data due to the inherent complexity. On the other hand, academic models either efficiently solve large-scale systems but lack accuracy (e.g., PyPSA [Brown et al., 2018], REMix [Gils et al., 2017]) or are tailored to small-scale systems (e.g., [Verheyen et al., 2024], [Lambert and Spliethoff, 2024], [Castelli et al., 2024]). Recent developments in the energy modeling community show an increase in integrating network design and operational decisions and apply this approach to micro-grids or short-term scenarios, while a solution for urban areas of large size and long-term scenarios is not available due to efficiency issues. Still, no solution effectively bridges the gap between these industrial and academic tools.
@@ 2 A Multi-Energy Integrated Operation and Design Optimization Problem
We introduce a novel network-flow-based model class integrating operation and design optimization for multi-energy systems. The operation optimization is modeled as a unit commitment problem [Sheble and Fahd, 1994], incorporating a time horizon, a set of generating and storage units, a representation of the topology, and demand/price forecast time series as input. Unit commitment involves three decisions:
@@ 2.1.3 Full Program:
The main goal is to fulfill the given demand while balancing all resources, represented by Constraint 1. Constraint 2 describes the conversion from one resource to another at a generating unit in form of a piecewise linear characteristic curve. Constraints 3 - 5 incorporate technical constraints such as activation, minimum up and down times, and ramping for generating units. Constraint 6 models the time linking of storage levels. Constraint 7 links investment decisions to the corresponding operation variables. Capacities are set for all non-binary variables by Constraints 8. For an in-depth discussion on the model, we refer to [Clarner et al., 2022].
@@ 3.1 Instance
[…]
@@ 3.2 Method
The goal is to explore various transformation pathways to achieve decarbonization targets while maintaining reasonable economic costs. Due to the instance size, finding all Pareto optimal solutions (i.e., solutions that cannot be improved in one objective without worsening another) for two objectives (costs, CO2 emissions) is intractable. Therefore, we explore a reasonable number of Pareto optimal solutions using a lexicographic optimization approach, prioritizing costs before CO2 emissions (hierarchical method [Marler and Arora, 2004]). While minimizing CO2 emissions, the optimal cost value is relaxed in increments of 5%, ranging up to 30% of the cost optimum. The 1% relaxation admits a 2% MIP-gap, the 5% relaxation a 1% MIP-gap, and further relaxations a 0.5% MIP-gap. Our computations were executed on a high-performance Cluster, featuring an Intel(R) Xeon(R) Gold 6342 CPU and comparable CPUs with a memory limit of 100 GB. Each individual solution was processed within a time frame of 1 to 6 hours.
@@ 3.3 Computational Results
11111.051.051.051.051.11.11.11.11.151.151.151.151.21.21.21.21.251.251.251.251.31.31.31.30.70.70.70.70.80.80.80.80.90.90.90.91111+ electrical heater- gas turbinenormalized costsnormalized $CO_{2}$ emissionsPareto optimal trade-offs

### v2

@@ Abstract
The need to decarbonize the energy system has intensified the focus on district heating networks in urban and suburban areas. Therefore, exploring transformation pathways with reasonable trade-offs between economic viability and environmental goals became necessary. We introduce a network-flow-based model class integrating unit commitment and long-term investment planning for multi-energy systems. While the integration of unit commitment and investment planning has been applied to multi-energy systems, a formal introduction and suitability for the application of long-term portfolio planning of an energy provider on an urban scale has yet to be met. Based on mixed integer linear programming, the model bridges the gap between overly detailed industrial modeling tools not designed for computational efficiency at scale and rather abstract academic models. The formulation is tested on Berlin’s district heating network. Hence, the challenge lies in a large number of variables and constraints and the coupling of time steps, for example, through investment decisions. A case study explores different solutions on the Pareto front defined by optimal trade-offs between minimizing costs and CO2 emissions through a lexicographic optimization approach. The resulting solution catalog can provide decision-makers valuable insights into feasible transformation pathways, highlighting distinctions between robust and target-dependent investments.The content of this report is also available as publication in Operations Research Proceedings 2024. Please always cite as: Riedmüller, S., Rivetta, F., Zittel, J. (2025). Long-Term Multi-objective Optimization for Integrated Unit Commitment and Investment Planning for District Heating Networks. In: Glomb, L. (eds) Operations Research Proceedings 2024. OR 2024. Lecture Notes in Operations Research. Springer, Cham. https://doi.org/10.1007/978-3-031-92575-7_33
Keywords:
multi-objective optimization multi-energy systems investment planning unit commitment
@@ 1 Introduction
While energy providers are urged to decarbonize, current optimization tools for multi-energy systems still fall into two categories: On one hand, detailed industrial modeling tools (e.g,. BoFiT by Volue [^1] https://www.volue.com/products-and-services/volue-bofit-optimisation-chp-sector-coupling-and-industry, ResOpt by Kisters [^2] https://www.kisters.eu/product/forecasting-and-optimisation/) do not allow for an integrated structural optimization for large input data due to the inherent complexity. On the other hand, academic models either efficiently solve large-scale systems but lack accuracy (e.g., PyPSA [1], REMix [4]) or are tailored to small-scale systems (e.g., [2], [8], [5]). Recent developments in the energy modeling community show an increase in integrating network design and operational decisions and apply this approach to micro-grids or short-term scenarios, while a solution for urban areas of large size and long-term scenarios is not available due to efficiency issues. Still, no solution effectively bridges the gap between these industrial and academic tools.
@@ 2 A Multi-Energy Integrated Operation and Design Optimization Problem
We introduce a novel network-flow-based model class integrating operation and design optimization for multi-energy systems. The operation optimization is modeled as a unit commitment problem [7], incorporating a time horizon, a set of generating and storage units, a representation of the topology, and demand/price forecast time series as input. Unit commitment involves three decisions:
@@ 2.1.3 Full Program:
The main goal is to fulfill the given demand while balancing all resources, represented by Constraint 1. Constraint 2 describes the conversion from one resource to another at a generating unit in form of a piecewise linear characteristic curve. Constraints 3 - 5 incorporate technical constraints such as activation, minimum up and down times, and ramping for generating units. Constraint 6 models the time linking of storage levels. Constraint 7 links investment decisions to the corresponding operation variables. Capacities are set for all non-binary variables by Constraints 8. For an in-depth discussion on the model, we refer to [3].
@@ 3.1 Instance
DemandBalanceHWE GasHWE GasCHP CoalDemandBalanceP2HCHP CoalCHP GasBalanceHWE $\geqDemandHWE GasDemandBalanceHWE GasCCGT$BalanceHWE GasCCGTHeatpumpHWE GasCHP GasDemandFuelFuelFuelPowerPowerPowerPowerPowerPower
@@ 3.2 Method
The goal is to explore various transformation pathways to achieve decarbonization targets while maintaining reasonable economic costs. Due to the instance size, finding all Pareto optimal solutions (i.e., solutions that cannot be improved in one objective without worsening another) for two objectives (costs, CO2 emissions) is intractable. Therefore, we explore a reasonable number of Pareto optimal solutions using a lexicographic optimization approach, prioritizing costs before CO2 emissions (hierarchical method [6]). While minimizing CO2 emissions, the optimal cost value is relaxed in increments of 5%, ranging up to 30% of the cost optimum. The 1% relaxation admits a 2% MIP-gap, the 5% relaxation a 1% MIP-gap, and further relaxations a 0.5% MIP-gap. Our computations were executed on a high-performance Cluster, featuring an Intel(R) Xeon(R) Gold 6342 CPU and comparable CPUs with a memory limit of 100 GB. Each individual solution was processed within a time frame of 1 to 6 hours.
@@ 3.3 Computational Results
111.051.051.11.11.151.151.21.21.251.251.31.30.70.70.80.80.90.911+ electrical heater- gas turbinenormalized costsnormalized $CO_{2}$ emissionsPareto optimal trade-offs
@@ 4.0.1 Acknowledgements
[…]
The authors have no competing interests to declare that are relevant to the content of this article.